9x+10=-8 find the value of x in given question EFQJNIFJEQAMOSLMALOJSOQIJOKCMLNVUFH

9x+10=-8
find the value of x in given question
EFQJNIFJEQAMOSLMALOJSOQIJOKCMLNVUFH
 

Grade:12th pass

2 Answers

Vikas TU
3 years ago
Dear student 
9x + 10 = -8 
9x = -8-10
9x = -18 
x = -18/9 
x = -2 
Hope this will help 
Good Luck 
Cheers ….............
Ram Kushwah
3 years ago
Given that
9x+10=-8
Taking 10 to RHS
9x=-8-10
9x=-18
Dividing both sides by 9
9x/9=-18/9
x=-2 Answer
Thanks
All the best